OpenStudy (anonymous):

could someone help me, give steps, on how to solve for x in an equation like this.... (in reply box) pleaase and thanks i really need help !

OpenStudy (anonymous):

\[4x+3 \over x-5\] is this a fraction? or is it mean to be divided..

OpenStudy (anonymous):

x-5>0 x>5

OpenStudy (gw2011):

This is a fraction and a fraction is also a division. If this is an equation, then I would set it equal to zero if it is not equal to anything else (that is, if this is a complete problem). Specifically, I would do the following: (4x+3)/(x-5) = 0 Multiply both sides of the equation by (x-5) 4x+3 = 0 4x = -3 x = -3/4

OpenStudy (anonymous):

it is a rational expression.

OpenStudy (gw2011):

A rational expression consists of rational numbers (integers) in which one can be written over another so long as the denominator is not zero. Therefore, x cannot equal 5 in the expression. The solution I arrived at is a rational expression. I hope this helps.